OpenStudy (anonymous):

Solve the equation. 42x – 5 = 64

OpenStudy (anonymous):

@demetra_harrison add 5 to both sides and show what u will get?

OpenStudy (anonymous):

okay thank

OpenStudy (anonymous):

42x – 5 = 64 42x = 64+5 42x = 69 x = 69/42

OpenStudy (anonymous):

64+5=69 69/42=1.64 x=1.64